# Learning Strategy Representation for Imitation Learning in Multi-Agent Games

### v1

@@ Introduction
Although reinforcement learning has become a powerful technique for sequential decision-making in various domains such as robotic manipulation [1], autonomous driving [7], and game playing [32], conventional reinforcement learning demands substantial online interactions with the environment, which can be costly and sample inefficient while potentially leading to safety risks [4, 5]. To address these issues, many methods have emerged to enable efficient learning using offline datasets generated by demonstrators. For example, imitation learning (IL) [26] replicates actions from the offline dataset without reward information, while offline reinforcement learning [13, 20] is provided access to reward signals. Offline learning datasets are usually collected from multiple demonstrators to enlarge dataset scale and diversity [30, 23, 24], which leads to a dataset of behaviors with various characteristics. However, standard IL algorithms treat all data samples in the dataset as homogeneous, potentially learning undesired behaviors from sub-optimal trajectories.
@@
  • We enhance existing IL methods by filtering out sub-optimal trajectories according to the RI and EL indicators, ensuring that IL is trained exclusively on the dominant trajectory.
@@
Imitation Learning. In the conventional IL settings, the expert trajectories only have information of state-action pairs without reward information [26, 27, 18, 10, 14], and it is assumed that the demonstrations are homogeneous oracle. However, realistic crowd-sourced datasets are usually multi-modal and include sub-optimal demonstrations. Some IL methods are proposed for multi-modal offline datasets, such as [16] and [11]. As for sub-optimal data, there are plenty of approaches to alleviate the negative influence [6, 8, 35], but all these methods require environment dynamics or the rankings over the demonstrations. In contrast, our method does not require such information. Sasaki and Yamashina [28] enhances behavior cloning (BC) with noisy demonstrations, but their method does not deal with general sub-optimal trajectories. TRAIL [33] achieves sample-efficient IL via a learned latent action space and a factored transition model. We would like to additionally mention the work by Franzmeyer et al. [12], which adopts a similar framework of filtering the offline dataset and uses an IL algorithm. Nevertheless, their method is studied in a cooperative setting and requires reward information.
IL with Representation Learning. The work by Beliaev et al. [3] closely aligns with our research, sharing the primary goal of extracting expertise levels of trajectories. They assume that the demonstrator has a vector indicating the expertise of latent skills, with each skill requiring a different level at a specific state. These elements jointly derive the expertise level. The method also considers the policy worse when it is closer to uniformly random distribution. However, this assumption cannot be satisfied even in simple games like RPS, where a uniformly random strategy constitutes a Nash equilibrium. Play-LMP [22] leverages unsupervised representation learning in a latent plan space for improved task generalization. However, employing a variational auto-encoder (VAE) with the encoder outputting latent plans is unsuitable for multi-player games, potentially leaking opponent information from the observations and disrupting the evaluation of the demonstrator. Grover et al. [15] also studies learning policy representations from offline trajectories. They use the information of agent identification during training, which enables them to add a loss to distinguish one agent from others. However, this information may not be provided in the offline datasets.
@@
Markov Games. A Markov game [21] is a partially observable Markov decision process [19] (POMDP) adapted to a multi-agent setting, where each agent has its own reward function. In a Markov game, there is a state space $S$ and $n$ agents, with each agent $i$ having a corresponding action space $A_{i}$ and observation space $O_{i}$. When an agent is not required to take action at a certain state, its action space contains only one action, referred to as a ‘null’ action. At each time step $t$: (1) each agent $i$ obtains an observation $o_{t}^{i}\in O_{i}$ and selects an action $a_{t}^{i}\in A_{i}$ based on the policy of agent $i$, $\pi_{i}:O_{i}\times A_{i}\to[0,1]$; (2) the agent receives a reward $\mathfrak{r}_{t}^{i}:S\times A_{i}\to\mathbb{R}$ based on the state and the action; (3) the state is changed according to the transition function $T:S\times A_{1}\times...\times A_{n}\to S$. For a complete trajectory $\tau=((o_{0}^{i},a_{0}^{i}),...,(o_{T}^{i},a_{T}^{i}))$ of agent $i$, there is a reward $\mathfrak{r}_{t}^{i}$ received at each time step $t$. We define the trajectory reward $\tau$ as $\hat{r}_{i}(\tau)=\sum_{t=0}^{T}\mathfrak{r}_{t}^{i}$. The expected trajectory reward of player $i$ with strategy $\pi_{i}$ and opponents with strategy $\pi_{-i}$ is defined as $r_{i}(\pi_{-i},\pi_{i})=\mathbb{E}_{\tau\sim(\pi_{-i},\pi_{i})}\left[\hat{r}_{i}(\tau)\right]$, in which $\tau\sim(\pi_{-i},\pi_{i})$ denote that $\tau$ is generated with agents using strategy $(\pi_{-i},\pi_{i})$.
@@
Best Response, Exploitability, and Nash Equilibrium. We use $r_{i}(\pi_{-i},\pi_{i})$ to specify the player strategy $\pi_{i}$ and opponent strategy $\pi_{-i}$. The best response of opponent strategy $\pi_{-i}$ is defined as $BR(\pi_{-i})=\text{argmax}_{\pi_{i}^{\prime}}r_{i}(\pi_{-i},\pi_{i}^{\prime})$, which refers to the strategy of player $i$ that maximizes player $i$’s reward. We additionally define the best response of strategy $\pi_{i}$ as $BR(\pi_{i})=\text{argmax}_{\pi_{-i}^{\prime}}\sum_{j\in P,j\neq i}r_{j}(\pi_{-i}^{\prime},\pi_{i})$, which equals to $\text{argmin}_{\pi_{-i}^{\prime}}r_{i}(\pi_{-i}^{\prime},\pi_{i})$ in the zero-sum case. $BR(\pi_{i})$ refers to the strategy of all the other players except player $i$ that maximizes their trajectory reward, which is equivalent to minimizing player $i$’s payoff in zero-sum games. Let $P$ be the set of all the players in the game, and the strategy $\pi=(\pi_{i})_{i\in P}$ be the strategy of all the players. We define the exploitability of strategy $\pi$ as $E(\pi)=\sum_{i\in P}\left(r_{i}(\pi_{-i},BR(\pi_{-i}))-r_{i}(\pi)\right)$, which reflects the extent to which the strategy can be exploited. In zero-sum symmetric cases, we define the exploitability of a player strategy $\pi_{i}$ as $E(\pi_{i})=-r_{i}(BR(\pi_{i}),\pi_{i})=\sum_{j\in P,j\neq i}r_{j}(BR(\pi_{i}),\pi_{i})$. A strategy $\pi_{i}$ is $\varepsilon$-Nash equilibrium if $E(\pi_{i})\leq\varepsilon$.
@@
Similar to [9], the loss function of P-VRNN is a negative of the variational lower bound, using Equations 1 and 5, as follows:
@@
It is the entropy of $p_{\theta}(a_{t}|z_{\leq t},a_{<t},o_{\leq t},l)$, which reflects the randomness of the player with strategy representation $l$, given $z_{\leq t},a_{<t}$, and $o_{\leq t}$. Following the hypothesis of Beliaev et al. [3], a strategy with more randomness is considered worse. Since we have the whole trajectory with a unified strategy representation $l$, we can define the RI of a trajectory as its cumulative reconstruction loss:
@@
where $r(\pi,\pi(\tau))$ returns the expected trajectory reward of a player with strategy $\pi(\tau)$ by default, $(x)^{+}=\max{\{x,0\}}$, and $\mathds{1}_{c}=1$ if and only if condition $c$ is satisfied, otherwise $\mathds{1}_{c}=0$. $EL(\tau)$ is the negative of the expectation of the trajectory reward less than 00 when played with the demonstrators who generate the offline dataset. This value can reflect the extent to which the demonstrators exploit the strategy of $\tau$. The reason why EL is chosen as an indicator is discussed in detail in the Appendix. To estimate EL with latent representation space structure, we provide an alternative definition of $EL_{\delta}$:
@@
where $d$ is a metric over the strategy representation space. Due to the Lipschitz continuity of the P-VRNN, the trajectories with similar strategy representations have similar strategies. Thus, to approximate the negative EL of $\tau$, we can calculate the mean of all the negative rewards of the trajectories with the strategy representations in the small neighborhood of $\tau$’s representation. It can be proved that $\lim_{\delta\to 0^{+}}EL_{\delta}(\tau)=EL(\tau)$. $EL_{\delta}$ has favorable properties, such as the low value near Nash equilibrium strategies. Given a trajectory $\tau$ and its corresponding distribution $\tau(\pi)$ over $\Pi$, $\pi(\tau)$ is $\epsilon_{1}$-Nash equilibrium, and we assume that any pure strategy can exploit another strategy by at most $M$. By the smoothness of $f$, we also assume that if $d(f(\tau_{1}),f(\tau_{2}))<\delta$, then $\int_{\pi\in\Pi}|\tau_{1}(\pi)-\tau_{2}(\pi)|\text{d}\pi<\alpha\delta$, where $\alpha$ is a constant. We can prove that $EL_{\delta}(\tau)<\epsilon_{1}+\alpha\delta M$ in the Appendix.
@@
We validate the effectiveness of our approach using two-player zero-sum games, including Two-player Pong, Limit Texas Hold’em [34], and Connect Four [31]. Details of each environment are provided in the Appendix.
(a) Player ID
(b) Randomness indicator
(c) Exploited level
(d) Trajectory reward
(e) Player ID
(f) Randomness indicator
(g) Exploited level
(h) Trajectory reward
(i) Player ID
(j) Randomness Indicator
(k) Exploited Level
(l) Trajectory reward
@@
Dataset generation. We employ different methods to create training datasets with diverse demonstrators for the environments. For Two-player Pong and Connect Four, we use self-play with opponent sampling [2] with the Proximal Policy Optimization (PPO) algorithm [29]. For Limit Texas Hold’em, we use neural fictitious self-play [17] with Deep Q-network (DQN) algorithm [25] to generate expert policies, given its complexity and the need to adapt to various opponents. Behavior models are then selected from multiple intermediate checkpoints to generate the offline data. We assume that only 5% of the dataset is reward-labeled for EL estimation. Details on the generation of demonstrators, their performance, and the dataset are provided in the Appendix.
@@
In this subsection, we visualize the learned strategy representations using multiple labels. If the latent space exceeds two dimensions, it is initially reduced to two dimensions using PCA. These reduced representations are then color-coded based on different labels: player ID, RI, EL, and trajectory reward. Note that player ID and trajectory reward serve as ground truth references while RI and EL are estimated. Instead of using the exact values, we color the percentiles of RI, EL, and trajectory reward. The implementation details for P-VRNN are provided in the Appendix.
@@
Table 1: $\operatorname{WS}$ of each IL algorithm over the demonstrator set, $\mathcal{I}$. Each algorithm was trained with distinct datasets filtered by various methods: (1) Original: utilizing the full dataset for training; (2) RI: filtering the dataset using the RI indicator; (3) EL: filtering the dataset using the EL indicator; and (4) Best: employing only the data generated by the dominant demonstrator, which serves as an oracle method. Higher is better.
@@
Two-player Pong. As shown in Figure 3a, strategy representations of each demonstrator naturally cluster together in the Two-player Pong environment. Figure 3d demonstrates that trajectory rewards only partially align with player strategies due to performance variability depending on the opponent’s strategy, a common characteristic of competitive games. Players $1$ and $8$, exhibiting the worst and best performances, respectively, form clusters with consistent values independent of the opponent. Figure 3b illustrates that the RI highlights players $5$, $6$, and $8$ as excelling in reconstruction tasks. Additionally, Figure 3c shows that the dominant players, $4$ and $8$, have strategies that are least susceptible to exploitation, signifying more robust performance. Additionally, it is observed that the most expansive cluster with the lowest density has the highest EL, suggesting that the least trained strategy exhibits unstable behavior and is the most vulnerable one to exploitation. These indicators establish a strong standard for data filtering in subsequent IL applications from two different perspectives, both differentiating between dominant and dominated strategies.
Limit Texas Hold’em & Connect Four. In Limit Texas Hold’em, there are seven players: two experts, three mid-level players, and two novice players. Figure 3e demonstrates that the learned representations are well separated and ordered according to their expertise levels. Figure 3h illustrates that while trajectory rewards can effectively identify very poor strategies, they fail to consistently differentiate among more effective strategies, as the rewards vary across different opponents. However, Figures 3f and 3g show that our proposed indicators not only perfectly distinguish the dominant strategies but also rank them accurately. In Connect Four, Figure 3i shows dominant player strategies on the right and dominated player strategies on the left. In contrast to the trajectory rewards which are inconsistent within the same strategy, our RI and EL patterns show a strong capability to extract characteristics and assess the performance of these strategies.
@@
To evaluate the STRIL, we considered three IL algorithms. First, we employed Behavior Cloning (BC), a basic IL algorithm. Next, we used IQ-Learn [14], an advanced imitative algorithm. Finally, we implemented ILEED [3], a state-of-the-art method capable of handling a diverse range of demonstrator data. In our evaluation, we excluded methods that rely on online interactions (e.g., GAIL [18]) or necessitate interactions with experts (e.g., DAgger [27]) in offline learning approaches. We applied STRIL to each algorithm to evaluate its performance enhancement. Note that all the experiments were repeated three times.
General results. In Table 1, we compared the WS of four types of data filtering methods. A hyperparameter search was conducted to identify the appropriate percentile, $p$, of indicators for each model and environment. Note that all experiments were repeated three times, and the results are reported with error bars. The original ILEED, which considers the expertise level of the data, generally performs better than other original algorithms on average. For the filtering method, the RI and EL enhance the performance of the original methods in most cases. In some instances, their performance is even comparable to the Best method. In the case of Two-player Pong, the EL method outperforms the RI method because EL more accurately distinguishes the dominant strategy. Additionally, in Limit Texas Hold’em, both RI and EL show similar performance, which aligns with the similar qualitative results observed in the strategy space. However, the RI and EL methods did not improve ILEED on Connect Four because the dataset aligns well with the assumption of ILEED. Consequently, filtering the data according to randomness is equivalent to reducing valid data, which results in worse performance. A detailed analysis of the dataset leading to this result is provided in the Appendix.

### v2

@@ Introduction
Although reinforcement learning has become a powerful technique for sequential decision-making in various domains such as robotic manipulation (Andrychowicz et al. 2020), autonomous driving (Chen, Yuan, and Tomizuka 2019), and game playing (Vinyals et al. 2019), conventional reinforcement learning demands substantial online interactions with the environment, which can be costly and sample inefficient while potentially leading to safety risks (Berner et al. 2019; Bojarski et al. 2016). To address these issues, many methods have emerged to enable efficient learning using offline datasets generated by demonstrators. For example, imitation learning (IL) (Pomerleau 1988) replicates actions from the offline dataset without reward information, while offline reinforcement learning (Fujimoto, Meger, and Precup 2019; Kumar et al. 2020) is provided access to reward signals. Offline learning datasets are usually collected from multiple demonstrators to enlarge dataset scale and diversity (Sharma et al. 2018; Mandlekar et al. 2019, 2021), which leads to a dataset of behaviors with various characteristics. However, standard IL algorithms treat all data samples in the dataset as homogeneous, potentially learning undesired behaviors from sub-optimal trajectories.
@@
  • We enhance existing IL methods by filtering out sub-optimal trajectories using the RI and EL indicators, ensuring that IL is trained only on the dominant trajectory.
@@
Imitation Learning. In the conventional IL settings, the expert trajectories only have information of state-action pairs without reward information (Pomerleau 1988; Ross, Gordon, and Bagnell 2011; Ho and Ermon 2016; Ding et al. 2019; Garg et al. 2021), and it is assumed that the demonstrations are homogeneous oracle. However, realistic crowd-sourced datasets are usually multi-modal and include sub-optimal demonstrations. Some IL methods are proposed for multi-modal offline datasets, such as (Hausman et al. 2017) and (Fei et al. 2020). As for sub-optimal data, there are plenty of approaches to alleviate the negative influence (Brown, Goo, and Niekum 2020; Chen, Paleja, and Gombolay 2021; Zhang et al. 2021; Kim et al. 2022; Xu et al. 2022), but all these methods require environment dynamics, the rankings over the demonstrations, or the identification of demonstrators. In contrast, our method does not require such information. Sasaki and Yamashina (2020) enhances behavior cloning (BC) with noisy demonstrations, but their method does not deal with general sub-optimal trajectories. TRAIL (Yang, Levine, and Nachum 2021) achieves sample-efficient IL via a learned latent action space and a factored transition model. We would like to additionally mention the work by Franzmeyer et al. (2024), which adopts a similar framework of filtering the offline dataset and uses an IL algorithm. Nevertheless, their method assumes a cooperative setting and requires reward information.
IL with Representation Learning. The work by Beliaev et al. (2022) closely aligns with our research, sharing the primary goal of extracting expertise levels of trajectories. They assume that the demonstrator has a vector indicating the expertise of latent skills, with each skill requiring a different level at a specific state. These elements jointly derive the expertise level. The method also considers the policy worse when it is closer to uniformly random distribution. However, this assumption cannot be satisfied even in simple games like RPS, where a uniformly random strategy constitutes a Nash equilibrium. Play-LMP (Lynch et al. 2020) leverages unsupervised representation learning in a latent plan space for improved task generalization. However, employing a variational auto-encoder (VAE) with the encoder outputting latent plans is unsuitable for multi-player games, potentially leaking opponent information from the observations and disrupting the evaluation of the demonstrator. Grover et al. (2018) also studies learning policy representations from offline trajectories. They use the information of agent identification during training, which enables them to add a loss to distinguish one agent from others. However, this information may not be provided in the offline datasets.
@@
Markov Games. A Markov game (Littman 1994) is a partially observable Markov decision process (Kaelbling, Littman, and Cassandra 1998) (POMDP) adapted to a multi-agent setting, where each agent has its own reward function. In a Markov game, there is a state space $S$ and $n$ agents, with each agent $i$ having a corresponding action space $A_{i}$ and observation space $O_{i}$. When an agent is not required to take action at a certain state, its action space contains only one action, referred to as a ‘null’ action. At each time step $t$: (1) each agent $i$ obtains an observation $o_{t}^{i}\in O_{i}$ and selects an action $a_{t}^{i}\in A_{i}$ based on the policy of agent $i$, $\pi_{i}:O_{i}\times A_{i}\to[0,1]$; (2) the agent receives a reward $\mathfrak{r}_{t}^{i}:S\times A_{i}\to\mathbb{R}$ based on the state and the action; (3) the state is changed according to the transition function $T:S\times A_{1}\times...\times A_{n}\to S$. For a complete trajectory $\tau=((o_{0}^{i},a_{0}^{i}),...,(o_{T}^{i},a_{T}^{i}))$ of agent $i$, there is a reward $\mathfrak{r}_{t}^{i}$ received at each time step $t$. We define the trajectory reward $\tau$ as $\hat{r}_{i}(\tau)=\sum_{t=0}^{T}\mathfrak{r}_{t}^{i}$. The expected trajectory reward of player $i$ with strategy $\pi_{i}$ and opponents with strategy $\pi_{-i}$ is defined as $r_{i}(\pi_{-i},\pi_{i})=\mathbb{E}_{\tau\sim(\pi_{-i},\pi_{i})}\left[\hat{r}_{i}(\tau)\right]$, in which $\tau\sim(\pi_{-i},\pi_{i})$ denote that $\tau$ is generated with agents using strategy $(\pi_{-i},\pi_{i})$.
@@
Best Response, Exploitability, and Nash Equilibrium. We use $r_{i}(\pi_{-i},\pi_{i})$ to specify the reward of the player playing $\pi_{i}$ against $\pi_{-i}$. The best response of opponent strategy $\pi_{-i}$ is defined as $BR(\pi_{-i})=\text{argmax}_{\pi_{i}^{\prime}}r_{i}(\pi_{-i},\pi_{i}^{\prime})$, which refers to the strategy of player $i$ that maximizes player $i$’s reward. We additionally define the best response of strategy $\pi_{i}$ as $BR(\pi_{i})=\text{argmax}_{\pi_{-i}^{\prime}}\sum_{j\in P,j\neq i}r_{j}(\pi_{-i}^{\prime},\pi_{i})$, which equals to $\text{argmin}_{\pi_{-i}^{\prime}}r_{i}(\pi_{-i}^{\prime},\pi_{i})$ in the zero-sum case. $BR(\pi_{i})$ refers to the strategy of all the other players except player $i$ that maximizes their trajectory reward, which is equivalent to minimizing player $i$’s payoff in zero-sum games. Let $P$ be the set of all the players in the game, and the strategy $\pi=(\pi_{i})_{i\in P}$ be the strategy of all the players. We define the exploitability of strategy $\pi$ as $E(\pi)=\sum_{i\in P}\left(r_{i}(\pi_{-i},BR(\pi_{-i}))-r_{i}(\pi_{-i},\pi_{i})\right)$, which reflects the extent to which the strategy can be exploited. In zero-sum symmetric cases, we define the exploitability of a player strategy $\pi_{i}$ as $E(\pi_{i})=-r_{i}(BR(\pi_{i}),\pi_{i})=\sum_{j\in P,j\neq i}r_{j}(BR(\pi_{i}),\pi_{i})$. A strategy $\pi_{i}$ is $\varepsilon$-Nash equilibrium if $E(\pi_{i})\leq\varepsilon$.
@@
Similar to (Chung et al. 2015), the loss function of P-VRNN is a negative of the variational lower bound, using Equations 1 and 5, as follows:
@@
It is the entropy of $p_{\theta}(a_{t}|z_{\leq t},a_{<t},o_{\leq t},l)$, which reflects the randomness of the player with strategy representation $l$, given $z_{\leq t},a_{<t}$, and $o_{\leq t}$. Following the hypothesis of Beliaev et al. (2022), a strategy with more randomness is considered worse. Since we have the whole trajectory with a unified strategy representation $l$, we can define the RI of a trajectory as its cumulative reconstruction loss:
@@
where $r(\pi,\pi(\tau))$ returns the expected trajectory reward of a player with strategy $\pi(\tau)$ by default, $(x)^{+}=\max{\{x,0\}}$, and $\mathds{1}_{c}=1$ if and only if condition $c$ is satisfied, otherwise $\mathds{1}_{c}=0$. $EL(\tau)$ is the negative of the expectation of the trajectory reward less than 00 when played with the demonstrators who generate the offline dataset. This value can reflect the extent to which the demonstrators exploit the strategy of $\tau$. To estimate EL with latent representation space structure, we provide an alternative definition of $EL_{\delta}$:
@@
where $d$ is a metric over the strategy representation space. Due to the Lipschitz continuity of the P-VRNN with respect to the representation, the trajectories with similar strategy representations have similar strategies. Thus, to approximate the negative EL of $\tau$, we can calculate the mean of all the negative rewards of the trajectories with the strategy representations in the small neighborhood of $\tau$’s representation. It can be proved that $\lim_{\delta\to 0^{+}}EL_{\delta}(\tau)=EL(\tau)$. $EL_{\delta}$ has favorable properties, such as the low value near Nash equilibrium strategies. Given a trajectory $\tau$ and its corresponding distribution $\tau(\pi)$ over $\Pi$, $\pi(\tau)$ is $\epsilon_{1}$-Nash equilibrium, and we assume that any pure strategy can exploit another strategy by at most $M$. We also assume that similar representations induce similar strategies: if $d(f(\tau_{1}),f(\tau_{2}))<\delta$, then $\int_{\pi\in\Pi}|\tau_{1}(\pi)-\tau_{2}(\pi)|\text{d}\pi<\alpha\delta$, where $\alpha$ is a constant. It can be proved that $EL_{\delta}(\tau)<\epsilon_{1}+\alpha\delta M$.
@@
We validate our approach using two-player zero-sum games: Two-player Pong, Limit Texas Hold’em (Zha et al. 2020), and Connect Four (Terry et al. 2021).
@@
Dataset generation. We employ different methods to create training datasets with diverse demonstrators for the environments. For Two-player Pong and Connect Four, we use self-play with opponent sampling (Bansal et al. 2018) with the Proximal Policy Optimization (PPO) algorithm (Schulman et al. 2017). For Limit Texas Hold’em, we use neural fictitious self-play (Heinrich and Silver 2016) with Deep Q-network (DQN) algorithm (Mnih et al. 2013) to generate expert policies, given its complexity and the need to adapt to various opponents. Behavior models are then selected from multiple intermediate checkpoints to generate the offline data. We assume that only 5% of the dataset is reward-labeled for EL estimation.
@@
In this subsection, we visualize the learned strategy representations using multiple labels. If the latent space exceeds two dimensions, it is initially reduced to two dimensions using PCA. These reduced representations are then color-coded based on different labels: player ID, RI, EL, and trajectory reward. Note that player ID and trajectory reward serve as ground truth references while RI and EL are estimated. Instead of using the exact values, we color the percentiles of RI, EL, and trajectory reward.
@@
Table 1: $\operatorname{WS}$ of each IL algorithm over the demonstrator set, $\mathcal{I}$. Each algorithm was trained with distinct datasets filtered by various methods: (1) Original: utilizing the full dataset for training; (2) RI: filtering the dataset using the RI indicator; (3) EL: filtering the dataset using the EL indicator; and (4) Best: employing only the data generated by the dominant demonstrator, which serves as an oracle method. Bold highlights the best performance among Original, RI, and EL, while underline shows if the ’Best’ method achieves the highest performance overall. Higher is better.
@@
Two-player Pong. As shown in Figure 3(a), strategy representations of each demonstrator naturally cluster together in the Two-player Pong environment. Figure 3(d) demonstrates that trajectory rewards only partially align with player strategies due to performance variability depending on the opponent’s strategy, a common characteristic of competitive games. Players $1$ and $8$, exhibiting the worst and best performances, respectively, form clusters with consistent values independent of the opponent. Figure 3(b) illustrates that the RI highlights players $5$, $6$, and $8$ as excelling in reconstruction tasks. Additionally, Figure 3(c) shows that the dominant players, $4$ and $8$, have strategies that are least susceptible to exploitation, signifying more robust performance. Additionally, it is observed that the most expansive cluster with the lowest density has the highest EL, suggesting that the least trained strategy exhibits unstable behavior and is the most vulnerable one to exploitation. These indicators establish a strong standard for data filtering in subsequent IL applications from two different perspectives, both differentiating between dominant and dominated strategies.
Limit Texas Hold’em & Connect Four. In Limit Texas Hold’em, there are seven players: two experts, three mid-level players, and two novice players. Figure 3(e) demonstrates that the learned representations are well separated and ordered according to their expertise levels. Figure 3(h) illustrates that while trajectory rewards can effectively identify very poor strategies, they fail to consistently differentiate among more effective strategies, as the rewards vary across different opponents. However, Figures 3(f) and 3(g) show that our proposed indicators not only perfectly distinguish the dominant strategies but also rank them accurately. In Connect Four, Figure 3(i) shows dominant player strategies on the right and dominated player strategies on the left. In contrast to the trajectory rewards which are inconsistent within the same strategy, our RI and EL patterns show a strong capability to extract characteristics and assess the performance of these strategies.
@@
To evaluate the STRIL, we considered three IL algorithms. First, we employed BC, a basic IL algorithm. Next, we used IQ-Learn (Garg et al. 2021), an advanced imitative algorithm. Finally, we implemented ILEED (Beliaev et al. 2022), a state-of-the-art method capable of handling a diverse range of demonstrator data. In our evaluation, we excluded methods that rely on online interactions (e.g., GAIL (Ho and Ermon 2016)) or necessitate interactions with experts (e.g., DAgger (Ross, Gordon, and Bagnell 2011)) in offline learning approaches. We applied STRIL to each algorithm to evaluate its performance enhancement. Note that all the experiments were repeated three times.
General results. In Table 1, we compared the WS of four types of data filtering methods. A hyperparameter search was conducted to identify the appropriate percentile, $p$, of indicators for each model and environment. Note that all experiments were repeated three times, and the results are reported with error bars. The original ILEED, which considers the expertise level of the data, generally performs better than other original algorithms on average. For the filtering method, the RI and EL enhance the performance of the original methods in most cases. In some instances, their performance is even comparable to the Best method. In the case of Two-player Pong, the EL method outperforms the RI method because EL more accurately distinguishes the dominant strategy. Additionally, in Limit Texas Hold’em, both RI and EL show similar performance, which aligns with the similar qualitative results observed in the strategy space. However, the RI and EL methods did not improve ILEED on Connect Four because the dataset aligns well with the assumption of ILEED. Consequently, filtering the data according to randomness is equivalent to reducing valid data, which results in worse performance.